# Kalman–Bucy Filtering with Randomized Sensing:

### v1

@@ A-A Proof of Lemma 1
$\bm{\Sigma}_{\lambda}=\lambda\bm{\Sigma}_{1}+(1-\lambda)\bm{\Sigma}_{2}$
with $\lambda\in[0,1]$.
@@
$\bm{A}\bm{\Sigma}_{\lambda}+\bm{\Sigma}_{\lambda}\bm{A}^{\top}+\bm{Q}_{c}$
$=\lambda(\bm{A}\bm{\Sigma}_{1}+\bm{\Sigma}_{1}\bm{A}^{\top}+\bm{Q}_{c})+(1-\lambda)(\bm{A}\bm{\Sigma}_{2}+\bm{\Sigma}_{2}\bm{A}^{\top}+\bm{Q}_{c})$
@@
$\bm{\Sigma}_{\lambda}\bm{G}\bm{\Sigma}_{\lambda}$
$=(\lambda\bm{\Sigma}_{1}+(1-\lambda)\bm{\Sigma}_{2})\bm{G}(\lambda\bm{\Sigma}_{1}+(1-\lambda)\bm{\Sigma}_{2})$
$=\lambda^{2}\bm{\Sigma}_{1}\bm{G}\bm{\Sigma}_{1}+(1-\lambda)^{2}\bm{\Sigma}_{2}\bm{G}\bm{\Sigma}_{2}$
$\quad+\lambda(1-\lambda)(\bm{\Sigma}_{1}\bm{G}\bm{\Sigma}_{2}+\bm{\Sigma}_{2}\bm{G}\bm{\Sigma}_{1})$
$=\lambda\bm{\Sigma}_{1}\bm{G}\bm{\Sigma}_{1}+(1-\lambda)\bm{\Sigma}_{2}\bm{G}\bm{\Sigma}_{2}$
$\quad-\lambda(1-\lambda)(\bm{\Sigma}_{1}\bm{G}\bm{\Sigma}_{1}+\bm{\Sigma}_{2}\bm{G}\bm{\Sigma}_{2}-(\bm{\Sigma}_{1}\bm{G}\bm{\Sigma}_{2}+\bm{\Sigma}_{2}\bm{G}\bm{\Sigma}_{1}))$
$=\lambda\bm{\Sigma}_{1}\bm{G}\bm{\Sigma}_{1}+(1-\lambda)\bm{\Sigma}_{2}\bm{G}\bm{\Sigma}_{2}$
$\quad-\lambda(1-\lambda)(\bm{\Sigma}_{1}-\bm{\Sigma}_{2})\bm{G}(\bm{\Sigma}_{1}-\bm{\Sigma}_{2})$
@@
$\bm{\Sigma}_{\lambda}\bm{G}\bm{\Sigma}_{\lambda}$
$\leq\lambda\bm{\Sigma}_{1}\bm{G}\bm{\Sigma}_{1}+(1-\lambda)\bm{\Sigma}_{2}\bm{G}\bm{\Sigma}_{2}$
@@
$\mathcal{R}_{\bm{G}}(\lambda\bm{\Sigma}_{1}+(1-\lambda)\bm{\Sigma}_{2})$
$=\bm{A}\bm{\Sigma}_{\lambda}+\bm{\Sigma}_{\lambda}\bm{A}^{\top}+\bm{Q}_{c}-\bm{\Sigma}_{\lambda}\bm{G}\bm{\Sigma}_{\lambda}$
$\geq\lambda(\bm{A}\bm{\Sigma}_{1}+\bm{\Sigma}_{1}\bm{A}^{\top}+\bm{Q}_{c})+\lambda\bm{\Sigma}_{1}\bm{G}\bm{\Sigma}_{1}$
$\quad+(1-\lambda)(\bm{A}\bm{\Sigma}_{2}+\bm{\Sigma}_{2}A^{\top}+\bm{Q}_{c})+(1-\lambda)\bm{\Sigma}_{2}\bm{G}\bm{\Sigma}_{2}$
$=\lambda\mathcal{R}_{\bm{G}}(\bm{\Sigma}_{1})+(1-\lambda)\mathcal{R}_{\bm{G}}(\bm{\Sigma}_{2})$
@@ A-B Proof of Lemma 2
The proof is inspired by [26, Theorem 2.1].
$\mathcal{R}_{\bm{G}}(\bm{\Sigma}_{1})-\mathcal{R}_{\bm{G}}(\bm{\Sigma}_{2})$

### v2

@@ A-A Proof of Lemma 1
$\bm{\Sigma}_{\beta}=\beta\bm{\Sigma}_{1}+(1-\beta)\bm{\Sigma}_{2}$
with $\beta\in[0,1]$.
@@
$\bm{A}\bm{\Sigma}_{\beta}+\bm{\Sigma}_{\beta}\bm{A}^{\top}+\bm{Q}_{c}$
$=\beta(\bm{A}\bm{\Sigma}_{1}+\bm{\Sigma}_{1}\bm{A}^{\top}+\bm{Q}_{c})+(1-\beta)(\bm{A}\bm{\Sigma}_{2}+\bm{\Sigma}_{2}\bm{A}^{\top}+\bm{Q}_{c})$
@@
$\bm{\Sigma}_{\beta}\bm{G}\bm{\Sigma}_{\beta}$
$=(\beta\bm{\Sigma}_{1}+(1-\beta)\bm{\Sigma}_{2})\bm{G}(\beta\bm{\Sigma}_{1}+(1-\beta)\bm{\Sigma}_{2})$
$=\beta^{2}\bm{\Sigma}_{1}\bm{G}\bm{\Sigma}_{1}+(1-\beta)^{2}\bm{\Sigma}_{2}\bm{G}\bm{\Sigma}_{2}$
$\quad+\beta(1-\beta)(\bm{\Sigma}_{1}\bm{G}\bm{\Sigma}_{2}+\bm{\Sigma}_{2}\bm{G}\bm{\Sigma}_{1})$
$=\beta\bm{\Sigma}_{1}\bm{G}\bm{\Sigma}_{1}+(1-\beta)\bm{\Sigma}_{2}\bm{G}\bm{\Sigma}_{2}$
$\quad-\beta(1-\beta)(\bm{\Sigma}_{1}\bm{G}\bm{\Sigma}_{1}+\bm{\Sigma}_{2}\bm{G}\bm{\Sigma}_{2}-(\bm{\Sigma}_{1}\bm{G}\bm{\Sigma}_{2}+\bm{\Sigma}_{2}\bm{G}\bm{\Sigma}_{1}))$
$=\beta\bm{\Sigma}_{1}\bm{G}\bm{\Sigma}_{1}+(1-\beta)\bm{\Sigma}_{2}\bm{G}\bm{\Sigma}_{2}$
$\quad-\beta(1-\beta)(\bm{\Sigma}_{1}-\bm{\Sigma}_{2})\bm{G}(\bm{\Sigma}_{1}-\bm{\Sigma}_{2})$
@@
$\bm{\Sigma}_{\beta}\bm{G}\bm{\Sigma}_{\beta}$
$\leq\beta\bm{\Sigma}_{1}\bm{G}\bm{\Sigma}_{1}+(1-\beta)\bm{\Sigma}_{2}\bm{G}\bm{\Sigma}_{2}$
@@
$\mathcal{R}_{G}(\beta\bm{\Sigma}_{1}+(1-\beta)\bm{\Sigma}_{2})$
$=\bm{A}\bm{\Sigma}_{\beta}+\bm{\Sigma}_{\beta}\bm{A}^{\top}+\bm{Q}_{c}-\bm{\Sigma}_{\beta}\bm{G}\bm{\Sigma}_{\beta}$
$\geq\beta(\bm{A}\bm{\Sigma}_{1}+\bm{\Sigma}_{1}\bm{A}^{\top}+\bm{Q}_{c})+\beta\bm{\Sigma}_{1}\bm{G}\bm{\Sigma}_{1}$
$\quad+(1-\beta)(\bm{A}\bm{\Sigma}_{2}+\bm{\Sigma}_{2}A^{\top}+\bm{Q}_{c})+(1-\beta)\bm{\Sigma}_{2}\bm{G}\bm{\Sigma}_{2}$
$=\beta\mathcal{R}_{G}(\bm{\Sigma}_{1})+(1-\beta)\mathcal{R}_{G}(\bm{\Sigma}_{2})$
@@ A-B Proof of Lemma 2
The proof is inspired by [freiling1996generalized, Theorem 2.1].
$\mathcal{R}_{G}(\bm{\Sigma}_{1})-\mathcal{R}_{G}(\bm{\Sigma}_{2})$